[Brahms-l] Shift report 20050605 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 05 2005 - 15:58:08 EDT
Shifter leader was sjs
Shifters were brk


Beam off since noon.
Since June 3:
2.3 deg 1723B 0.6M(trig2)
         861B 1.6M
         861A 0.7M   (--finished during shift)
60 deg   700A 0.46M(trig3)
52.5 deg 700A 0.74M
40 deg   700A 0.49M   (--finished during shift)
D5 polarity has now been changed to B, MRS at 60 deg.

Problem with trigger 6 resolved...

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 5 15:58:21 2005

This archive was generated by hypermail 2.1.8 : Sun Jun 05 2005 - 15:58:34 EDT